OpenStudy (wwe123):

if α,β are the roots of the equation 2x^2 -3x+1 = 0 form the equation whose equation are ....

OpenStudy (wwe123):

sorry if α,β are the roots of the equation 2x^2 -3x+1 = 0 form the equation whose roots are

OpenStudy (wwe123):

\[\frac{ \alpha }{ 2\beta+3} , \frac{ \beta }{ 2\alpha+3 }\]

OpenStudy (anonymous):

so we need to find alpha, beta .. can you solve 2x^2 -3x +1 =0 ?

OpenStudy (anonymous):